Question 2

--/1

View Policies

Current Attempt in Progress

Logan Miller started her own accounting firm, Miller Accounting, on June 1, 2021. Logan Miller wants to prepare monthly financial statements, so adjusting journal entries are required on June 30. Selected transactions for June follow:

1. $980 of supplies were used during the month.
2. Utilities expense incurred but not yet recorded or paid on June 30, 2021, is $271.
3. Paid cash of $3,864 for a one-year insurance policy on June 1, 2021. The policy came into effect on this date.
4. On June 1, purchased office equipment for $14,520 cash. It is being depreciated at $242 per month for 60 months.
5. On June 1, Logan signs a note payable for $12,200, 6% interest.
6. June 30 is a Wednesday and employees are paid on Fridays. Miller Accounting has two employees, who are paid $800 each for a five-day workweek that ends on Friday.
7. On June 15, received a $1,200 advance cash payment from a client for accounting services expected to be provided in the future. As at June 30, one half of these services had not been performed.
8. Invoices representing $1,700 of services performed during the month of June have not been recorded as at June 30.


Prepare adjusting entries for the items above. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)
Answer #1
Adjusting Entries
No. Accounts Debit Credit
1 Supplies Expense Dr $980
To Supplies $980
(Being supplies used during the month.)
2 Utilities expense Dr $271
To Utility Expense Payable $271
(Being Utilities expense incurred but not yet recorded or paid.)
3 Insurance expense Dr $322 =3864/12*1
Prepaid Insurance Dr $3,542 =3864/12*11
To Cash $3,864
(Being paid cash of $3,864 for a one-year insurance policy on June 1, 2021.)
4 Depreciation Dr $242
To Office Equipment $242
(Being Depreciation of office equipment for June.)
5 Interest Expense Dr $61 =12200*6%/12
To Interest payable on Note $61
(Being interest @ 6% on $12200 note payable for June.)
6 Salaries Expense Dr $960 =800*2*3/5
To Salaries payable $960
(Being salary payable for 2 employees for 3 out of 5 days.)
7 Revenue Dr $600 =1200/2
To Unearned Revenue $600
(Being received a $1,200 advance cash payment from a client earlier,one half of these services had not been performed.)
8 Accounts Receivable Dr $1,700
To Revenue $1,700
(Being services performed during June, not yet recorded.)